OpenStudy (anonymous):

somebody can help me THERMODYNAMICS *a thermodynamic steady flow system receives 4.56 kg per min of a fluid where p1=137.90 kPa, v1=0.0388m^3/kg, w1=122m/s, and u1=17.16kJ/kg. the fluid leaves the system at a boundary where p2=551.6kPa, v2=0.193m^3/kg, w2=183m/s and u2=52.80 kJ/kg. during passage through the system the fluid receives 3000J/s of heat. determine the work.

OpenStudy (anonymous):

enthalpy: h1=u1+p1*v1=17.16+137.9*0.0388=22.51 KJ/Kg h2=u2+p2*v2=52.8+551.6*0.193=159.2588 KJ/Kg

OpenStudy (anonymous):

Energy Balance for Open Systems without accumulation \[Q-W+\sum m _{i}(h _{i}+\frac{w _{i}^2}{2}+gz _{i})-\sum m _{e}(h _{e}+\frac{w _{e}^2}{2}+gz _{e})=0\]

OpenStudy (anonymous):

Q=3000J/s=3 KJ/s m=4.56 Kg/min=0.076 Kg/s

OpenStudy (anonymous):

then \[W=3+0.076(22.51+\frac{122^2}{2000}-159.2588-\frac{183^2}{2000})\]

OpenStudy (anonymous):

W is in KJ/s or KW use a calculator to evaluate it

OpenStudy (anonymous):

W=-8.099 KJ/s or KW

OpenStudy (anonymous):

the answer was given and it was -486.0018KJ/s ...the difficult here is the conversion

OpenStudy (anonymous):

let me check it again

OpenStudy (anonymous):

my equation for W is rifht or not?

OpenStudy (anonymous):

*right

OpenStudy (anonymous):

there is a two work flow, given, so i think it is wf1=mp1v1

OpenStudy (anonymous):

thats work of flow i think the problem is calculating work of shaft

OpenStudy (anonymous):

@miat the answer is given maybe is in KJ/min because my answer*60=answer is given could u plz check the question again

OpenStudy (anonymous):

the w1& w2 by the way is velocity

OpenStudy (anonymous):

well thats right

OpenStudy (anonymous):

W=-8.099 KJ/s or KW is correct but in order to have the final answer used this formula: w=(KE1-KE2)+(Wf1-Wf2)+(V1-V2)+Q

OpenStudy (anonymous):

oh yes thats correct

OpenStudy (anonymous):

KE=m w^2/2 Wf=mPv V=mu so whats your problem with conversion?